prove ||Ax|| < ||A|| ||x||


1
Expert's answer
2022-03-09T14:29:55-0500

Solution


Let us suppose that


"||Ax|| > ||A|| ||x||"


"\Rightarrow\frac{||Ax||}{||x||} > ||A||"


"\Rightarrow|A\frac{x}{||x||}| > ||A||"


Since "\frac{x}{||x||}|" is a unit vector, therefore, from above we can write,


"\Rightarrow ||A|| > ||A||" which constradicts


Hence


"||Ax|| < ||A|| ||x||"


Proved